# Evaluating link prediction: New perspectives and recommendations

### v2

@@ 5.2 Distance between the end nodes in a test node pair
Here we aim to understand whether prediction performance of various LP methods vary when the distance between the two nodes in test node pairs differ. For each dataset, we performed paired t-test to determine if the prediction performance differ in terms of their absolute values for the two cases: two-hop away test node pairs and three-hop away test node pairs. Also, we measure the Kendall’s $\tau$ values for the same to estimate the ordinal association between the two cases. We consider AUROC, AUPR (balanced), AUPR (imbalanced), $Pr@P$ (balanced) and $Pr@P$ (imbalanced) as the evaluation measures. We performed the experiment across all datasets and prediction types.
Table 4 shows that the absolute performance of the prediction methods differs significantly in most of the cases across all datasets, prediction types and evaluation metrics. Overall, for a given dataset, the winner between two-hop and three-hop is consistent across the evaluation measures. As far as the missing LP is concerned, there is no clear winner between the two cases. However, for the future LP, three-hop seems to do better than its two-hop counterpart, where, three-hop wins in clg-msg-t, math-flow-t, ast-phys-t and lst-fm-t, but there is no clear winner for the rest three datasets. When we compare the performance for the networks (wherever applicable) based on prediction type, we find some inconsistencies, which is most prominent in the Ast-phys network, where its missing LP version is dominated by two-hop, but three-hop does better in its future LP counterpart. These results indicate that the LP methods produce different level of accuracy for predicting links based on the distance of the end nodes in the network. Moreover, this behavior differs over the prediction types. From the Kendall’s Tau results it is also evident that the rankings of the prediction methods for the two cases vary significantly. All correlation values can be found in Table 17 at Appendix C.
Overall, the obtained results support the need to evaluate LP methods separately for the two-hop and three-hop case to understand its behavior in distance controlled manner, and our experimental setup provides an opportunity to measure it. Moreover, as the local LP methods are applicable only on the two-hop case, but the global methods are applicable in both the cases, our experimental setup provides a robust environment to compare a local with a global method.
@@ 5.3 Prediction method type
Table 5 shows that the prediction performance of the prediction methods in the three groups differ significantly for almost all cases in two-hop. Post-hoc test for the missing LP in two-hop cases presented in Table 6 shows that across almost all datasets learning group is the clear winner, but although global-sim methods outperforms local-sim in a few datasets and evaluation methods, in most cases there are no significant difference. For the two-hop cases in future LP shown in Table 7, although in clg-msg-t, math-flow-t and math-flow-t datasets, learning methods perform best, results are not conclusive for the rest. Table 8 shows that, for three-hop case, there is no significant difference in performance between the global-sim and learning based methods, except for few cases like bio-yeast, forum, blog-cat and lst-fm-t. For bio-yeast, forum and blog-cat, global-sim methods win, and learning based methods win in lst-fm-t. However, the number of the test node pairs for bio-yeast, forum and blog-cat in three-hop case is low, so, over-fitting might have caused the performance degradation for the learning based methods.
Overall, the results indicate that the relative performance of different types of LP methods vary with the distance between the end nodes. Moreover, even when the distance is same, the relative performance of the method types vary with the prediction type. The results in this section advocate for the proposed distance controlled and prediction type controlled setup, along with the consideration of the type of the methods while comparing performance of multiple methods, in order to evaluate an LP method in a robust and detailed manner.
@@ 5.4 Network type
As most of the LP methods apply on undirected networks, while applying those methods on directed networks, they are converted to undirected ones ignoring directions. Here we investigate whether this information loss affects their prediction performance. For each of future and missing LP, we divide the datasets into two groups: directed and undirected, and perform Student’s t-test for a LP method given its accuracy against a particular evaluation metric. To avoid getting overloaded with the high number of the LP methods, we conduct the test once for each type of methods: local-sim, global-sim and learning, by considering the performance score of the best performing method in that group. In Table 9, we report the test results. For the best performing methods in each group given an evaluation metric, readers may refer to Appendix B.
@@ 5.5 Class imbalance and early retrieval
In this section, we analyze how class imbalance affects the evaluation of LP. First we investigate if the relative performance of the LP methods differ subject to AUROC vs. AUPR. Both of AUROC and AUPR consider all threshold values, but AUROC is not affected by class imbalance whereas AUPR is. We consider both balanced as well as imbalanced (skew level $1:10$) test set for AUPR. Next, we take the imbalance sensitive measure AUPR, and compare the relative performance of the LP methods on it for the balanced and imbalanced version of the test set. We further do the same with the single threshold imbalance sensitive metric $Pr@P$. We consider Kendall’s Tau Rank Correlation Coefficient ($\tau$) to measure the (in)consistency of ranks of the LP methods when evaluated in each pair of setups stated above. The $\tau$ values for each of the datasets are shown in Table 10. We summarize our observations as follows.
@@ 5.5.1 Early retrieval:
For certain applications like recommender systems, LP methods must perform well at the early retrieval phase, and it becomes critical when the data is highly skewed. Here we investigate how class imbalance affects early retrieval. The ROC curve, rather the Concentrated ROC (CROC) curve [34] can reflect the performance in the early retrieval stage. However, as it is skew insensitive, it is unable to reflect the performance in imbalanced setup. Saito et al. [32] has shown how PRC can be used to understand the effect of imbalance in early retrieval. We first investigate how the performance on the single point metric $Pr@P/2$ (both balanced and imbalanced versions) correlate with that of AUROC, and the balanced and imbalanced versions of $Pr@P/2$ correlate with each other, and then show an example ROC plot demonstrating the different behavior of the prediction methods in the two test settings. For each dataset, we measure the Kendall’s Tau Rank Correlation Coefficient ($\tau$) of the LP methods for the pair of cases stated above in our controlled setup.
@@
5.6 Summary of the observation and recommendations
Our main findings from the experiments are summarized as follows:
  • Future LP vs. Missing LP: LP methods perform better when future LP is considered than that of missing LP
  • Distance between the end nodes: behavior of the LP methods are not consistent on the two test sets: two-hop and three-hop away node pairs
  • Prediction method type: when absolute performance of the LP methods are compared based on their category, we found significant difference in performance for the test set consisting of two-hop away node pairs; but for its three-hop away counterpart, no significant difference was observed
  • Network type: converting directed networks to undirected ones and applying LP methods on it may significantly affect its performance of LP, specifically for future LP on the test set of two-hop away node pairs
  • Class imbalance: injecting imbalance in the test set affects the relative performance of the LP methods for the skew sensitive metrics like AUPR and $Pr@P$, and the rankings based on AUPR and $Pr@P$ are not consistent with that of the skew insensitive metric AUROC
  • Early retrieval: rankings of the LP methods differ based on their performance in the early retrieval phase when imbalanced test set used as opposed to balanced one.
Based on our findings, we present our recommendations to evaluate a LP method as following.
Recommendations Controlled experimental setup: A LP method should be evaluated for both the problem types: Missing LP and Future LP, on a carefully designed distance controlled and skew controlled test set, following the controlled experimental setup presented in Section 4. Intra-group and Inter-group comparison: A method’s performance needs to be compared within its category, as well as with all the methods irrespective of their categories. Evaluation measure: We recommend AUPR as a single point summary evaluation measure to be employed on both of the balanced and imbalanced test set. Along with it, PRC, on both of balanced and imbalanced test set, should be investigated to focus on certain portion of the whole spectrum of the thresholds, such as the early retrieval region, which may be necessary for certain applications. Another single point metric: $Pr@P/2$, can be useful to investigate performance in the early retrieval region.
6 Conclusion, Limitations, and Future Directions
@@
This paper considered evaluating link prediction on a simple, unweighted, undirected and homogeneous graph, and analyzed the LP methods which ignore time information. This work can be extended to identify the unexplored avenues in evaluation of LP on specialized graphs, such as weighted, directed and heterogeneous. Various aspects of the performance of temporal LP methods and LP methods that are specifically designed for the directed networks, which this paper did not consider, may be investigated as further studies.
@@ 7.2 Funding
Appendix A Description of Network Datasets
A.1 Networks with timestamps:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
A.2 Networks without timestamps:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
Appendix B Consolidated Results: Best Performing Methods
@@
Appendix C Kendall’s $\tau$ Results for Section 5.1 and Section 5.2
NETWORK AUROC AUPR Pr@P
balanced imbalanced balanced imbalanced
two-hop CollegeMsg 0.184 0.184 0.219 0.156 0.203
Email-Eu 0.272 0.319 0.472 0.294 0.526
Mathoverflow 0.339 0.386 0.414 0.326 0.488
Ast-phys 0.487 0.433 0.377 0.525 0.393
lastFm 0.477 0.465 0.381 0.462 0.335
IAcontact 0.027 0.080 0.038 0.039 0.075
FB-forum 0.321 0.347 0.438 0.311 0.271
Topology 0.785 0.783 0.771 0.723 0.742
Mooc 0.744 0.743 0.698 0.678 0.741
three-hop CollegeMsg 0.009 0.118 0.158 0.022 0.230
Email-Eu -0.030 0.085 0.167 -0.144 0.181
Mathoverflow -0.020 -0.071 0.018 0.139 0.060
Ast-phys 0.256 0.464 0.578 0.422 0.628
lastFm
IAcontact
FB-forum 0.189 0.208 0.283 0.203 0.334
Topology 0.749 0.747 0.738 0.649 0.773
Mooc
Table 17: Kendall’s $\tau$ correlation results for two-hop vs three-hop
DATASET AUROC AUPR Pr@P
balanced imbalanced balanced imbalanced
missing LP clg-msg 0.275 0.203 0.241 0.316 0.207
email -0.059 -0.012 -0.155 -0.091 0.007
math-flow 0.399 0.385 0.227 0.311 0.196
ast-phys 0.278 0.391 0.572 0.373 0.542
lst-fm
forum 0.184 0.227 0.236 0.221 0.221
topo 0.233 0.316 0.302 0.176 0.241
act-mooc
pow-grid -0.144 0.084 0.173 -0.108 0.189
routers 0.242 0.005 0.079 0.285 0.068
bio-yeast 0.216 0.260 0.362 0.316 0.302
fb -0.178 0.450 0.211 -0.159 0.025
blog-cat 0.213 0.031 0.049 0.319 0.033
future LP clg-msg-t 0.580 0.589 0.548 0.578 0.489
email-t 0.603 0.443 0.390 0.609 0.337
math-flow-t 0.387 0.437 0.467 0.391 0.445
ast-phys-t 0.660 0.705 0.677 0.508 0.653
lst-fm-t 0.112 -0.021 -0.095 0.222 -0.131
forum-t 0.476 0.467 0.511 0.488 0.406
topo-t 0.363 0.391 0.360 0.326 0.361
act-mooc-t 0.439 0.589 0.573 0.413 0.594

### v4

@@ 6.2 Distance between the end nodes in a test node pair
Here we aim to understand whether prediction performance of various LP methods vary when the distance between the two nodes in test node pairs differ. For each dataset, we performed paired t-test to determine if the prediction performance differ in terms of their absolute values for the two cases: two-hop away test node pairs and three-hop away test node pairs. Also, we measure the Kendall?s $\tau$ values for the the same to estimate the ordinal association between the two cases. We consider AUROC, AUPR (balanced), AUPR (imbalanced), $Pr@P$ (balanced) and $Pr@P$ (imbalanced) as the evaluation measures. We performed the experiment across all datasets and prediction types. We summarize the results of the two tests in Table 5 and Table 6 respectively.
Table 5 shows that the absolute performance of the prediction methods differs significantly in most of the cases across all datasets, prediction types and evaluation metrics. Overall, for a given dataset, the winner between two-hop and three-hop is consistent across the evaluation measures. As far as the missing LP is concerned, there is no clear winner between the two cases. However, for the future LP, three-hop seems to do better than its two-hop counterpart, where, three-hop wins in clg-msg-t, math-flow-t, ast-phys-t and lst-fm-t, but there is no clear winner for the rest three datasets. When we compare the performance for the networks (wherever applicable) based on prediction type, we find some inconsistencies, which is most prominent in the Ast-phys network, where its missing LP version is dominated by two-hop, but three-hop does better in its future LP counterpart. These results indicate that the LP methods produce different level of accuracy for predicting links based on the distance of the end nodes in the network. Moreover, this behavior differs over the prediction types. From Table 6 it is also evident that the rankings of the prediction methods for the two cases vary significantly. Overall, the obtained results support the need to evaluate LP methods separately for the two-hop and three-hop case to understand its behavior in distance controlled manner, and our experimental setup provides an opportunity to measure it. Moreover, as the local LP methods are applicable only on in the two-hop case, but the global methods are applicable in both the cases, our experimental setup provides a robust environment to compare a local with a global method.
@@
Table 6: Kendall’s $\tau$ correlation results for two-hop vs three-hop
DATASET AUROC AUPR Pr@P
balanced imbalanced balanced imbalanced
missing LP clg-msg 0.275 0.203 0.241 0.316 0.207
email -0.059 -0.012 -0.155 -0.091 0.007
math-flow 0.399 0.385 0.227 0.311 0.196
ast-phys 0.278 0.391 0.572 0.373 0.542
lst-fm
forum 0.184 0.227 0.236 0.221 0.221
topo 0.233 0.316 0.302 0.176 0.241
act-mooc
pow-grid -0.144 0.084 0.173 -0.108 0.189
routers 0.242 0.005 0.079 0.285 0.068
bio-yeast 0.216 0.260 0.362 0.316 0.302
fb -0.178 0.450 0.211 -0.159 0.025
blog-cat 0.213 0.031 0.049 0.319 0.033
future LP clg-msg-t 0.580 0.589 0.548 0.578 0.489
email-t 0.603 0.443 0.390 0.609 0.337
math-flow-t 0.387 0.437 0.467 0.391 0.445
ast-phys-t 0.660 0.705 0.677 0.508 0.653
lst-fm-t 0.112 -0.021 -0.095 0.222 -0.131
forum-t 0.476 0.467 0.511 0.488 0.406
topo-t 0.363 0.391 0.360 0.326 0.361
act-mooc-t 0.439 0.589 0.573 0.413 0.594
@@ 6.3 Prediction method type
Table 7 shows that the prediction performance of the prediction methods in the three groups differ significantly for almost all cases in two-hop. Post-hoc test for the missing LP in two-hop cases presented in Table 8 shows that across almost all datasets learning group is the clear winner, but although global-sim methods outperforms local-sim in a few datasets and evaluation methods, in most cases there are no significant difference. For the two-hop cases in future LP shown in Table 9, although in clg-msg-t, math-flow-t and math-flow-t datasets, learning methods perform best, results are not conclusive for the rest. Table 10 shows that, for three-hop case, there is no significant difference in performance between the global-sim and learning based methods, except for few cases like bio-yeast, forum, blog-cat and lst-fm-t. For bio-yeast, forum and blog-cat, global-sim methods win, and learning based methods win in lst-fm-t. However, the number of the test node pairs for bio-yeast, forum and blog-cat in three-hop case is low, so, over-fitting might have caused the performance degradation for the learning based methods. Overall, the results indicate that the relative performance of different types of LP methods vary with the distance between the end nodes. Moreover, even when the distance is same, the relative performance of the method types vary with the prediction type. The results in this section advocate for the proposed distance controlled and prediction type controlled setup, along with the consideration of the type of the methods while comparing performance of multiple methods, in order to evaluate an LP method in a robust and detailed manner.
@@ 6.4 Network type
As most of the LP methods apply on undirected networks, while applying those methods on directed networks, they are converted to undirected ones ignoring directions. Here we investigate whether this information loss affects their prediction performance. For each of future and missing LP, we divide the datasets into two groups: directed and undirected, and perform Student’s t-test for a LP method given its accuracy against a particular evaluation metric. To avoid getting overloaded with the high number of the LP methods, we conduct the test once for each type of methods: local-sim, global-sim and learning, by considering the performance score of the best performing method in that group. In Table 11, we report the test results. For the best performing methods in each group given an evaluation metric, readers may refer to Appendix A.
@@ 6.5 Class imbalance and Early retrieval
In this section, we analyze how class imbalance affects the evaluation of LP. First we investigate if the relative performance of the LP methods differ subject to AUROC vs. AUPR. Both of AUROC and AUPR consider all threshold values, but AUROC is not affected by class imbalance whereas AUPR is. We consider both balanced as well as imbalanced (skew level $1:10$) test set for AUPR. Next, we take the imbalance sensitive measure AUPR, and compare the relative performance of the LP methods on it for the balanced and imbalanced version of the test set. We further do the same with the single threshold imbalance sensitive metric $Pr@P$. We consider Kendall?s Tau Rank Correlation Coefficient ($\tau$) to measure the (in)consistently of ranks of the LP methods when evaluated in each pair of setups stated above. The $\tau$ values for each of the datasets are shown in Table 12. We summarize our observations as follows.
@@ 6.5.1 Early Retrieval:
For certain applications like recommender systems, LP methods must perform well at the early retrieval phase, and it becomes critical when the data is highly skewed. Here we investigate how class imbalance affects early retrieval. The ROC curve, rather the Concentrated ROC (CROC) curve [34] can reflect the performance in the early retrieval stage. However, as it is skew insensitive, it is unable to reflect the performance in imbalanced setup. Saito et al. [32] has shown how PRC can be used to understand the effect of imbalance in early retrieval. We first investigate how the performance on the single point metric $Pr@P/2$ (both balanced and imbalanced versions) correlate with that of AUROC, and the balanced and imbalanced versions of $Pr@P/2$ correlate with each other, and then show an example ROC plot demonstrating the different behavior of the prediction methods in the two test settings. For each dataset, we measure the Kendall?s Tau Rank Correlation Coefficient ($\tau$) of the LP methods for the pair of cases stated above in our controlled setup.
@@
6.6 Summary
Our proposed experimental setup allowed us to conduct the tests presented in this section in order to evaluate the existing LP methods in a controlled environment. Our main findings are: (a) LP methods perform better when future LP is considered than that of missing LP; (b) behavior of the LP methods are not consistent on the two test sets: two-hop and three-hop away node pairs; (c) when absolute performance of the LP methods are compared based on their category, we found significant difference in performance for the test set consisting of two-hop away node pairs, but for its three-hop away counterpart, no significant difference was observed; (d) converting directed networks in undirected ones and applying LP methods on it may significantly affect its performance for future LP on the test set of two-hop away node pairs; (e) injecting imbalance in the test set affects the relative performance of the LP methods for the skew sensitive metrics like AUPR and $Pr@P$, and the rankings based on AUPR and $Pr@P$ are not consistent with that of the skew insensitive metric AUROC; (f) rankings of the LP methods differ based on their performance in the early retrieval phase when imbalanced test set used as opposed to balanced one.
Based on our findings, we recommend evaluating the LP methods in the controlled experimental setup presented in this paper. The methods should be evaluated on both the problem types, on the carefully designed distance controlled and skew controlled test set. A method’s performance needs to be compared within its category in this setup, as well as with all the methods irrespective of their categories. We recommend AUPR as a single point summary evaluation measure to be employed on both of the balanced and imbalanced test set. Along with it, PRC, on both of balanced and imbalanced test set, should be investigated to focus on certain portion of the whole spectrum of the thresholds, such as the early retrieval region, which may be necessary for certain applications.
7 Conclusion and Future Directions
@@
This paper considered evaluating link prediction on a simple, unweighted, undirected and homogeneous graph, and analyzed the LP methods which ignore time information. This work can be extended to identify the unexplored avenues in evaluation of LP on specialized graphs, such as weighted, directed and heterogeneous. Various aspects of the performance of temporal LP methods and LP methods, which are specifically designed for the directed networks, may be investigated as further studies.
@@ 8.2 Funding
Appendix A Consolidated results: Best performing methods in individual groups of methods